# On measures derived from orbital integrals

### v2

@@ Abstract.
In this article we propose a novel framework to derive piecewise polynomial measures which result from invariant measures on adjoint orbits in the general context of compact and semi-simple Lie groups. The measures are computed from orbital integrals by a series of transformations on spaces of polynomials endowed with the apolar inner product.
Keywords. orbital integral, orbital measures, piecewise polynomial measure, apolar inner product, Fisher Bombieri inner product, reproducing kernel.
@@ 1. Introduction
In this article we develop a novel approach to the derivation of some piecewise polynomial measures or Duistermat-Heckman type measures from Harish-Chandras (HC) orbital integral formula in the context of compact semi-simple Lie groups. These measures are derived from orbital measures, that is, invariant probability measures on adjoint orbits of compact semi-simple Lie groups. Specifically, we derive the pushforward of an orbital measure by the projection to the Cartan algebra, and we derive the radial part of the convolution of two orbital measures. These problems were previously addressed in [Z16, O13, F15] and in [F19, CMZ19, Z18]. They are probabilistic versions of the Schur-Horn theorem [H54] and Horn’s problem [H62]. Given hermitian matrices $X$ and $Y$ with fixed eigenvalues the Schur-Horn theorem characterizes the possible diagonal entries of $X$ as the convex closure of the permutation of eigenvalues, while Horn’s problem asks for a characterization of the possible eigenvalues of the sum $X+Y$. The possible eigenvalues are given by a set of linear inequalities and the problem was solved by Klyatchko in [K98].
@@
The article is organized as follows. In Section 2 we present the necessary results to state and prove the main theorems of the article. In Section 3 we give a characterization of the pushforward of an orbital measure and in Section 4 we give a characterization of the radial part of the convolution of two orbital measures.
@@ 2. Preliminaries
We review the results which are necessary to prove the main results of the article, such as the Harish-Chandra integral formula and operators between spaces of polynomials endowed with the apolar inner product. Most results are known, to the best of our knowledge only Propositions 2.5, 2.7, 2.9 and 2.10 are new.
@@ 2.2. Spaces of polynomials and operators between them
The creation and annihilation operator are fundamental in the theory of Segal-Bargman space. They consist of multiplication by a variable and differentiation by the same variable. In the context of spaces of polynomials we define for $\alpha\in V$ the multiplication operator
$M_{\alpha}:f(x)\operatorname{\mathcal{P}}^{k}(V)\to\langle\alpha,x\rangle f(x)\operatorname{\mathcal{P}}^{k}(V),\qquad g(x)\mapsto\langle\alpha,x\rangle g(x)$
and the derivative operator
$A_{\alpha}:\langle\alpha,x\rangle f(x)\operatorname{\mathcal{P}}^{k}(V)\to f(x)\operatorname{\mathcal{P}}^{k}(V),\qquad g(x)\mapsto\langle\alpha,\partial\rangle g(x).$
Proposition 2.6.
For $k\in\operatorname{\mathbb{N}}_{0}$ and $\alpha\in V$
$M_{\alpha}^{*}=A_{\alpha}.$
Proof.
For $g$ in the domain of $M_{\alpha}$ and $h$ in the domain of $A_{\alpha}$ we have
$\left[M_{\alpha}g,h\right]$ $=g(\partial)\langle\alpha,\partial\rangle h(x)|_{x=0}$
$=g(\partial)A_{\alpha}h(x)|_{x=0}$
$=\left[g,A_{\alpha}h\right],$
where we used the definition of the apolar inner product. ∎
The inverse of the annihilation and creation operators are the division operator $D_{\alpha}=M_{\alpha}^{-1}$ and the anti-derivative operator $I_{\alpha}=A_{\alpha}^{-1}$. They are adaptations of the adjoints of the divided difference or Bernstein-Gelfand-Gelfand operators introduced in Section $6$ of [PS09]. For $k\in\operatorname{\mathbb{N}}_{0}$, $\alpha\in V$ and a polynomial $f\in\operatorname{\mathcal{P}}(V)$ we define the division operator
$D_{\alpha}:\langle x,\alpha\rangle f(x)\operatorname{\mathcal{P}}^{k}(V)\to f(x)\operatorname{\mathcal{P}}^{k}(V)$
as
$g(x)\mapsto D_{\alpha}g(x)=\frac{1}{\langle\alpha,x\rangle}g(x),$
and the anti-derivative operator
$I_{\alpha}:f(x)\operatorname{\mathcal{P}}^{k}(V)\to\langle x,\alpha\rangle f(x)\operatorname{\mathcal{P}}^{k}(V)$
as
$g(x)\mapsto I_{\alpha}g(x)=\int^{\langle x,\alpha^{\circ}\rangle}_{0}g(x-\alpha t)dt,$
where $\alpha^{\circ}=\alpha/\langle\alpha,\alpha\rangle$. Note that the integration is done on the segment from $x$ to the projection of $x$ to the hyperplane $\{\alpha\}^{\perp}$. For $\alpha$ such that $\|\alpha\|=1$, the signed distance of $x$ to the hyperplane $\{\alpha\}^{\perp}$ is $\langle x,\alpha\rangle$. The proof of the following proposition is similar to the proof of Lemma $6.3$ in [PS09].
Proposition 2.7.
For the operators defined above we have
$D_{\alpha}^{*}=I_{\alpha},$
that is, the adjoint of the division operator is the anti-derivative.
Proof.
For a constant $c\neq 0$ and unit norm $\alpha\in V$ we have $D_{c\alpha}=\frac{1}{c}D_{\alpha}$. Also, by a linear change of variables in the definition of $I_{\alpha}$ we can assume that $\|\alpha\|=1$. Since an orthogonal change of coordinates leaves the equation unchanged we can prove the identity for the standard basis vector $\alpha=e_{1}$. We can also verify the identity for polynomials $x_{1}^{k}f(x_{2},\dots,x_{n})$ and $x_{1}^{l}g(x_{2},\dots,x_{n})$ since polynomials of this type span the spaces of polynomials. Since the operators $D_{e_{1}}$ and $I_{e_{1}}$ don’t act on the factors $f$ and $g$ and since
$\left[x_{1}^{k}f(x_{2},\dots,x_{n}),x_{1}^{l}g(x_{2},\dots,x_{n})\right]=\left[x_{1}^{k},x_{1}^{l}\right]\left[f(x_{2},\dots,x_{n}),g(x_{2},\dots,x_{n})\right]$
it is enough to verify the identity in the monomials $x_{1}^{k}$ and $x_{1}^{l}$. We have
$D_{e_{1}}(x_{1}^{k})=\frac{1}{x_{1}}x_{1}^{k}=x_{1}^{k-1}$
and
$I_{e_{1}}(x_{1}^{l})=\int_{0}^{x_{1}}t^{l}dt=\frac{1}{l+1}x_{1}^{l+1}.$
It is easy to very that $\left[x_{1}^{k-1},x_{1}^{l}\right]=\left[x_{1}^{k},\frac{1}{l+1}x_{1}^{l+1}\right]$. If $k\neq l+1$ then both side are zero and if $k=l+1$ then both sides are equal to $l!$. ∎
Until the end of this section we consider vector spaces which are the Cartan algebras $\operatorname{\mathfrak{t}}$ of Lie algebras of compact semi-simple Lie groups. One of the essential properties of the discriminant $\Delta:\operatorname{\mathfrak{t}}\to\operatorname{\mathbb{R}}$ is that it skew with respect to the action of $\operatorname{\mathcal{W}}$, $\Delta(w(x))=\operatorname{\epsilon}(w)\Delta(x)$ for $w\in\operatorname{\mathcal{W}}$. This follows from the fact that if $\alpha$ is a simple root the reflection through the plane $\langle\alpha,x\rangle=0$ sends $\alpha\mapsto-\alpha$ and permutes the other positive roots. The next proposition is well known.
@@
For each reflection $w_{\alpha}\in\operatorname{\mathcal{W}}$ we have $f(w_{\alpha}(x))=-f(x)$ so $f$ vanishes on the hyperplane $\langle\alpha,x\rangle=0$ and the linear polynomial $\langle\alpha,x\rangle$ divides $f$. Since the Lie algebra is semi-simple its root system is reduced, that is, the only scalar multiple of a root $\alpha$ is $\alpha$. Therefore, the positive roots are relatively prime, since they only divide each other if they are scalar multiples. Since for each positive root the polynomial $\langle\alpha,x\rangle=0$ divides $f$ the discriminant $\Delta$ divides the polynomial $f$. ∎
@@
Therefore we can define the division by the discriminant
$D_{\Delta}:\operatorname{\mathcal{P}}^{k+|\Phi^{+}|}_{\operatorname{alt}}(\operatorname{\mathfrak{t}})\to\operatorname{\mathcal{P}}^{k}_{\operatorname{sym}}(\operatorname{\mathfrak{t}})\qquad\mbox{given by}\qquad D_{\Delta}f(x)=\frac{f(x)}{\Delta(x)}.$
We can define $D_{\Delta}=D_{\alpha_{l}}\circ D_{\alpha_{l-1}}\circ\dots\circ D_{\alpha_{2}}\circ D_{\alpha_{1}}$ for an ordering $\alpha_{1},\alpha_{2},\dots,\alpha_{l}$ of the $l=|\Phi^{+}|$ positive roots. The order of the composition
$\Delta(x)\operatorname{\mathcal{P}}^{k}_{\operatorname{sym}}(\operatorname{\mathfrak{t}})\xrightarrow{D_{\alpha_{1}}}\prod_{\alpha\neq\alpha_{1}}\langle\alpha,x\rangle\operatorname{\mathcal{P}}^{k}_{\operatorname{sym}}(\operatorname{\mathfrak{t}})\xrightarrow{D_{\alpha_{2}}}\prod_{\alpha\neq\alpha_{1},\alpha_{2}}\langle\alpha,x\rangle\operatorname{\mathcal{P}}^{k}_{\operatorname{sym}}(\operatorname{\mathfrak{t}})\xrightarrow{D_{\alpha_{3}}}\ldots$
$\ldots\xrightarrow{D_{\alpha_{l-2}}}\langle\alpha_{l-1},x\rangle\langle\alpha_{l},x\rangle\operatorname{\mathcal{P}}^{k}_{\operatorname{sym}}(\operatorname{\mathfrak{t}})\xrightarrow{D_{\alpha_{l-1}}}\langle\alpha_{l},x\rangle\operatorname{\mathcal{P}}^{k}_{\operatorname{sym}}(\operatorname{\mathfrak{t}})\xrightarrow{D_{\alpha_{l}}}\operatorname{\mathcal{P}}^{k}_{\operatorname{sym}}(\operatorname{\mathfrak{t}})$
does not change the final outcome. The operator
$I_{\Delta}:\operatorname{\mathcal{P}}^{k}_{\operatorname{sym}}(\operatorname{\mathfrak{t}})\to\operatorname{\mathcal{P}}^{k+|\Phi^{+}|}_{\operatorname{alt}}$
is defined by $I_{\Delta}=I_{\alpha_{1}}\circ I_{\alpha_{2}}\circ\dots\circ I_{\alpha_{l-1}}\circ I_{\alpha_{l}}$ for an ordering $\alpha_{1},\alpha_{2},\dots\alpha_{l}$ of the $l=|\Phi^{+}|$ positive roots $\Phi^{+}$.
@@
The operator $I_{\Delta}=I_{\alpha_{1}}\circ I_{\alpha_{2}}\circ\dots\circ I_{\alpha_{l-1}}\circ I_{\alpha_{l}}$ is independent of the ordering of positive roots and it satisfies
$I_{\Delta}=D_{\Delta}^{*}.$
@@
Since the order of the composition $D_{\Delta}=D_{\alpha_{l}}\circ D_{\alpha_{l-1}}\circ\dots\circ D_{\alpha_{2}}\circ D_{\alpha_{1}}$ is irrelevant and since $I_{\alpha}=D_{\alpha}^{*}$ for all $\alpha\in\operatorname{\mathfrak{t}}$ by Proposition 2.7 the conclusion of the statement of the theorem follows. ∎
We denote with $\mathbf{1}$ the constant polynomial $\mathbf{1}(x)=1$.
@@
The iterated anti-derivative $I_{\Delta}$ satisfies
@@
We have $D_{\Delta}(\operatorname{\mathbb{R}}\Delta)=\operatorname{\mathbb{R}}\mathbf{1}$ and $D_{\Delta}(\{\operatorname{\mathbb{R}}\Delta\}^{\perp})=\{\operatorname{\mathbb{R}}\mathbf{1}\}^{\perp}$. Therefore $I_{\Delta}\mathbf{1}=D^{*}_{\Delta}\mathbf{1}=d\Delta$ for a $d\in\operatorname{\mathbb{R}}$. We have
$1=\left[\mathbf{1},\mathbf{1}\right]=\left[D_{\Delta}\Delta,\mathbf{1}\right]=\left[\Delta,I_{\Delta}\mathbf{1}\right]=\left[\Delta,d\Delta\right]=d\left[\Delta,\Delta\right].$

### v3

@@ Abstract.
In this article we propose a novel framework to derive piecewise polynomial measures which result from invariant measures on adjoint orbits in the general context of compact and semi-simple Lie groups. The measures are computed from orbital integrals by a series of transformations on spaces of polynomials endowed with the apolar inner product. In the case of the unitary group we give a formula for the moments of the projection of an orbital measure.
Keywords. orbital integral, orbital measures, piecewise polynomial measure, moments of measure, apolar inner product, Fisher Bombieri inner product, reproducing kernel.
@@ 1. Introduction
In this article we develop a novel approach to the derivation of some piecewise polynomial measures or Duistermat-Heckman type measures from Harish-Chandras (HC) orbital integral formula in the context of compact semi-simple Lie groups. These measures are derived from orbital measures, that is, invariant probability measures on adjoint orbits of compact semi-simple Lie groups. Specifically, we derive the pushforward of an orbital measure by the projection to the Cartan algebra, and we derive the radial part of the convolution of two orbital measures. In the case of type A root system we give a formula for the moments of the projection of an orbital measure. The characterization of these measures were previously addressed in [Z16, O13, F15] and in [F19, CMZ19, Z18]. They are probabilistic versions of the Schur-Horn theorem [H54] and Horn’s problem [H62]. Given hermitian matrices $X$ and $Y$ with fixed eigenvalues the Schur-Horn theorem characterizes the possible diagonal entries of $X$ as the convex closure of the permutation of eigenvalues, while Horn’s problem asks for a characterization of the possible eigenvalues of the sum $X+Y$. The possible eigenvalues are given by a set of linear inequalities and the problem was solved by Klyatchko in [K98].
@@
The article is organized as follows. In Section 2 we present the necessary results to state and prove the main theorems of the article. In Section 3 we give a characterization of the pushforward of an orbital measure, and we find the moments of the measure for type A root systems. In Section 4 we give a characterization of the radial part of the convolution of two orbital measures.
@@ 2. Preliminaries
We review the results which are necessary to prove the main results of the article, such as the Harish-Chandra integral formula and operators between spaces of polynomials endowed with the apolar inner product. Most results are known, to the best of our knowledge Propositions 2.5, Proposition 2.9 and Proposition 2.14 are new.
@@ 2.2. Spaces of polynomials and operators between them
Until the end of this section we consider vector spaces which are the Cartan algebras $\operatorname{\mathfrak{t}}$ of Lie algebras of compact semi-simple Lie groups. One of the essential properties of the discriminant $\Delta:\operatorname{\mathfrak{t}}\to\operatorname{\mathbb{R}}$ is that it skew with respect to the action of $\operatorname{\mathcal{W}}$, $\Delta(w(x))=\operatorname{\epsilon}(w)\Delta(x)$. This follows from the fact that if $\alpha$ is a simple root the reflection through the plane $\langle\alpha,x\rangle=0$ sends $\alpha\mapsto-\alpha$ and permutes the other positive roots. The next proposition is well known.
@@
For each reflection $w_{\alpha}\in\operatorname{\mathcal{W}}$ we have $f(w_{\alpha}(x))=-f(x)$ so $f$ vanishes on the hyperplane $\langle\alpha,x\rangle=0$ and the polynomial $\langle\alpha,x\rangle$ divides $f$. Since the Lie algebra is semi-simple its root system is reduced, that is, the only scalar multiple of a root $\alpha$ is $\alpha$. Therefore, the positive roots are relatively prime, since they only divide each other if they are scalar multiples. Since for each positive root $\langle\alpha,x\rangle=0$ divides $f$ the discriminant $\Delta$ divides the polynomial $f$
@@
Definition 2.7.
The division by the discriminant is defined as
$D_{\Delta}:\operatorname{\mathcal{P}}^{k+|\Phi^{+}|}_{\operatorname{alt}}(\operatorname{\mathfrak{t}})\to\operatorname{\mathcal{P}}^{k}_{\operatorname{sym}}(\operatorname{\mathfrak{t}})\qquad\mbox{given by}\qquad D_{\Delta}f(x)=\frac{f(x)}{\Delta(x)}$
and its adjoint $I_{\Delta}=D_{\Delta}^{*}$ is
$I_{\Delta}=D_{\Delta}^{*}:\operatorname{\mathcal{P}}^{k}_{\operatorname{sym}}(\operatorname{\mathfrak{t}})\to\operatorname{\mathcal{P}}^{k+|\Phi^{+}|}_{\operatorname{alt}}(\operatorname{\mathfrak{t}}).$
For the case $\operatorname{\mathfrak{t}}\simeq\operatorname{\mathbb{R}}$ the polynomials are of one variable and we get the characterization of $D_{\Delta}^{*}$ as an anti-derivative $I_{\Delta}f(x)=\int_{0}^{x}f(t)dt$.
@@
In the case of the one dimensional Cartan algebra $\operatorname{\mathfrak{t}}\simeq\operatorname{\mathbb{R}}$ the adjoint $D_{\Delta}^{*}=I_{\Delta}$ of the division operator is the anti-derivative.
@@
It is enough to verify the identity in the monomials $x^{k}$ and $x^{l}$. We have
$D_{\Delta}(x^{k})=\frac{1}{x}x^{k}=x^{k-1}$
and
$I_{\Delta}(x^{l})=\int_{0}^{x}t^{l}dt=\frac{1}{l+1}x^{l+1}.$
It is easy to very that $\left[x^{k-1},x^{l}\right]=\left[x^{k},\frac{1}{l+1}x^{l+1}\right]$. If $k\neq l+1$ then both side are zero and if $k=l+1$ then both sides are equal to $l!$. ∎
The adjoint $D_{\Delta}^{*}$ of $D_{\Delta}$ can be characterized as follows in the general case
@@
The operator $D_{\Delta}^{*}$ is the inverse of the discriminant with differential variables $\Delta(\partial)$, that is,
$D_{\Delta}^{*}=\Delta(\partial)^{-1}.$
Proof.
For symmetric polynomials $f$ and $g$
$\left[f,g\right]$ $=\left[D_{\Delta}(\Delta.f),g\right]$
$=\left[\Delta.f,D_{\Delta}^{*}g\right]$
$=\left[f,\Delta(\partial)D_{\Delta}^{*}g\right],$
therefore $\Delta(\partial)D_{\Delta}^{*}=Id$. Also, for alternating polynomials $r$ and $s$
$\left[r,s\right]$ $=\left[\Delta.D_{\Delta}(r),s\right]$
$=\left[D_{\Delta}(r),\Delta(\partial)s\right]$
$=\left[r,D_{\Delta}^{*}\Delta(\partial)s\right],$
so that $D_{\Delta}^{*}\Delta(\partial)=Id$. ∎
We have also the following characterization of $I_{\Delta}$ on the functions $q^{k}_{a}$.
Proposition 2.10.
For $k\in\operatorname{\mathbb{N}}_{0}$ and $a\in\operatorname{\mathfrak{t}}$ such that $\Delta(a)\neq 0$ the identity
$I_{\Delta}(q_{a}^{k})=\frac{1}{\Delta(a)}r_{a}^{k+|\Phi^{+}|}$
holds.
Proof.
Using the fact that $\langle\partial,\alpha\rangle e^{\langle x,a\rangle}=\langle\alpha,a\rangle e^{\langle x,a\rangle}$ we obtain the following well known identity for power series
$\Delta(\partial)\left(\frac{1}{|\mathcal{W}|}\sum_{w\in\mathcal{W}}\epsilon(w)e^{\langle x,w(a)\rangle}\right)=\Delta(a)\sum_{w\in\mathcal{W}}e^{\langle x,w(a)\rangle}.$
If we take polynomial parts by applying the truncation operator $F^{k}$ we get
$\Delta(\partial)\left(r_{a}^{k+|\Phi^{+}|}\right)=\Delta(a)q_{a}^{k}.$
If we apply $\Delta(\partial)^{-1}$ in the case $\Delta(a)\neq 0$ we obtain the formula stated in the theorem. ∎
Next we give a computation in the case of polynomials of lowest degree. We denote with $\mathbf{1}$ the constant polynomial $\mathbf{1}(x)=1$.
Proposition 2.11.
The operator $I_{\Delta}$ satisfies
@@
We have $D_{\Delta}(\operatorname{\mathbb{R}}\Delta)=\operatorname{\mathbb{R}}\mathbf{1}$ and $D_{\Delta}(\{\operatorname{\mathbb{R}}\Delta\}^{\perp})=\{\operatorname{\mathbb{R}}\mathbf{1}\}^{\perp}$. Therefore $I_{\Delta}\mathbf{1}=D^{*}_{\Delta}\mathbf{1}=d\Delta$ for $d\in\operatorname{\mathbb{R}}$. We have
$1=\left[D_{\Delta}\Delta,\mathbf{1}\right]=\left[\Delta,I_{\Delta}\mathbf{1}\right]=\left[\Delta,d\Delta\right]=d\left[\Delta,\Delta\right].$
@@
In the case of type A root systems one can give the following characterization of the operator $I_{\Delta}=D_{\Delta}^{*}$ by computing its matrix coefficients and taking the transpose of a matrix. We assume that the group is the unitary group $\operatorname{U}(n)$, its center is not trivial but this makes the link to the theory of symmetric polynomials more direct. The unitary orbital integral in this case is known as the Harish-Chandra-Itzykson-Zuber (HCIZ) integral [IZ80]. It is usually written
$\int_{\operatorname{U}(n)}e^{\operatorname{Tr}(uAu^{-1}B)}du$ $=\left(\prod_{p=1}^{n-1}p!\right)\frac{\operatorname{det}{\left[e^{a_{i}b_{j}}\right]_{i,j=1}^{n}}}{\Delta(A)\Delta(B)}$
$=\left(\prod_{p=1}^{n}p!\right)\frac{\frac{1}{n!}\sum_{w\in S_{n}}e^{\langle A,w(B)\rangle}}{\Delta(A)\Delta(B)}$
where $\operatorname{det}$ is the determinant of a matrix, $\operatorname{U}(n)$ is the group of $n$-by-$n$ unitary matrices, $A$ and $B$ are fixed $n$-by-$n$ diagonal matrices with eigenvalues $a_{1}<\dots<a_{n}$ and $b_{1}<\dots<b_{n}$ respectively, and
$\Delta(A)=\prod_{i<j}(a_{j}-a_{i})$
is the Vandermonde determinant.
Remark 2.12.
The HCIZ has become an important identity in quantum field theory, random matrix theory, and algebraic combinatorics.
Let $\operatorname{\mathbb{N}}_{0}$ stand for the non negative integers. For $\mu\in\operatorname{\mathbb{N}}_{0}^{n}$ we denote the monomials as usual with $x^{\mu}=x_{1}^{\mu_{1}}\dots x_{n}^{\mu_{n}}$ and we use the notations $\mu!=\mu_{1}!\dots\mu_{n}!$ and $|\mu|=\mu_{1}+\dots+\mu_{n}$. The set $\Pi$ of partitions is defined as
$\Pi=\{\lambda\in\operatorname{\mathbb{N}}_{0}^{n}:\lambda_{1}\geq\lambda_{2}\geq\dots\geq\lambda_{n}\}.$
For $k\in\operatorname{\mathbb{N}}_{0}$ we denote with
$\Pi^{k}=\{\lambda\in\Pi:|\lambda|\leq k\}$
the partitions of total size not greater than $k$. We set $\delta=(n-1,n-2,\dots,0)$. For $\lambda\in\Pi$, if $(x_{1},\dots,x_{n})$ are the eigenvalues of a hermitian matrix $x$ we define
$\chi_{\lambda}(x)=s_{\lambda}(x_{1},\dots,x_{n}),$
where $s_{\lambda}$ is a Schur polynomial. These polynomials are defined by
$s_{\lambda}(x_{1},\dots,x_{n})=\frac{a_{\lambda+\delta}(x_{1},\dots,x_{n})}{a_{\delta}(x_{1},\dots,x_{n})},$
where
$a_{\mu}(x_{1},\dots,x_{n})=\operatorname{det}\left[x_{i}^{\mu_{j}}\right]_{i,j=1}^{n}.$
Note that $a_{\delta}(x_{1},\dots,x_{n})=\Delta(x_{1},\dots,x_{n})$ is the Vandermonde determinant.
Remark 2.13.
The irreducible polynomial representations of the general linear group are labelled by Young diagrams, which we think of as vectors $\lambda\in\Pi$. The character of the $\lambda$-representation is given by $\chi_{\lambda}$, the Schur polynomial evaluated at the eigenvalues of an invertible matrix.
For $\mu\in\Pi$ we denote with $m_{\mu}$ the monomial symmetric polynomial which we defined as the sum of $x^{\lambda}$ where $\lambda$ ranges over distinct permutations of $\mu$, therefore $m_{\mu}$ is the sum of $|\mathcal{W\cdot\mu}|=|S_{n}\cdot\mu|$ distinct monomials. Here $|S_{n}\cdot\mu|$ is the cardinality of the Weyl orbit of $\mu$.
Schur polynomials can be expressed as linear combinations of monomial symmetric functions with non-negative integer coefficients $K_{\lambda\mu}$ called Kostka numbers
$s_{\lambda}=\sum_{\mu}K_{\lambda\mu}m_{\mu}.$
The Kostka numbers $K_{\lambda\mu}$ are given by the number of semi-standard Young tableaux of shape $\lambda\in\Pi$ and weight $\mu\in\Pi$. References for Kostka numbers are Section VI.1 in [M05] and Theorem 2.11.2 in [S01] for the representation theoretic aspect.
It is known that $(m_{\mu})_{\mu\in\Pi^{k}}$ and $(a_{\lambda+\delta})_{\delta\in\Pi^{k}}$ form algebraic bases of $\operatorname{\mathcal{P}}^{k}_{\operatorname{sym}}(\operatorname{\mathbb{R}}^{n})$ and $\operatorname{\mathcal{P}}^{k+l}_{\operatorname{alt}}(\operatorname{\mathbb{R}}^{n})$ respectively, where $l=|\Phi^{+}|=\frac{n(n-1)}{2}$. We have
$\left[a_{\lambda+\delta},a_{\lambda+\delta}\right]=n!(\lambda+\delta)!\qquad\mbox{and}\qquad\left[m_{\mu},m_{\mu}\right]=|S_{n}\cdot\mu|\mu!.$
We scale these polynomials to be of unit norm and define for $\lambda\in\Pi^{k}$
$a^{\prime}_{\lambda+\delta}=\frac{1}{\sqrt{n!(\lambda+\delta)}}a_{\lambda+\delta}\qquad\mbox{and}\qquad m^{\prime}_{\mu}=\frac{1}{|S_{n}\cdot\mu|\mu!}m_{\mu}.$
Therefore, for $k\in\operatorname{\mathbb{N}}$ the set $(a^{\prime}_{\lambda+\delta})_{\lambda\in\Pi^{k}}$ is an orthonormal basis of $\operatorname{\mathcal{P}}_{\operatorname{alt}}^{k+l}(\operatorname{\mathbb{R}}^{n})$ and the set $(m^{\prime}_{\mu})_{\mu\in\Pi^{k}}$ is an orthonormal basis of $\operatorname{\mathcal{P}}_{\operatorname{sym}}^{k}(\operatorname{\mathbb{R}}^{n})$.
We have the following characterization of the operator $I_{\Delta}$ given by its matrix coefficients.
Proposition 2.14.
For $k\in\operatorname{\mathbb{N}}$ the operator $D_{\Delta}:\operatorname{\mathcal{P}}^{k+l}_{\operatorname{alt}}(\operatorname{\mathbb{R}}^{n})\to\operatorname{\mathcal{P}}^{k}_{\operatorname{sym}}(\operatorname{\mathbb{R}}^{n})$ and its adjoint $I_{\Delta}=D_{\Delta}^{*}$ are given by
$D_{\Delta}a_{\lambda+\delta}=\sum_{\lambda\in\Pi}K_{\lambda\mu}m_{\mu}$
and
$I_{\Delta}m_{\mu}=D^{*}_{\Delta}m_{\mu}=\sum_{\lambda\in\Pi}\frac{|S_{n}\cdot\mu|\mu!}{n!(\lambda+\delta)!}K_{\lambda\mu}a_{\lambda+\delta}.$
Proof.
Note that for $\lambda\in\Pi^{k}$
$D_{\Delta}a_{\lambda+\delta}=\frac{a_{\lambda+\delta}}{a_{\delta}}=s_{\lambda}=\sum_{\lambda}K_{\lambda\mu}m_{\mu},$
Therefore
$D_{\Delta}a^{\prime}_{\lambda+\delta}=s_{\lambda}=\sum_{\lambda}K^{\prime}_{\lambda\mu}m^{\prime}_{\mu},$
with
$K^{\prime}_{\lambda\mu}=\frac{\sqrt{|S_{n}\cdot\mu|\mu!}}{\sqrt{n!(\lambda+\delta)!}}K_{\lambda\mu}.$
Hence for $\mu\in\Pi^{k}$
$D^{*}_{\Delta}m^{\prime}_{\mu}=\sum_{\lambda}K^{\prime}_{\lambda\mu}a^{\prime}_{\lambda+\mu},$
and this implies that
$D^{*}_{\Delta}m_{\mu}=\sum_{\lambda}\frac{|S_{n}\cdot\mu|\mu!}{n!(\lambda+\delta)!}K_{\lambda\mu}a_{\lambda+\delta}.$